OpenStudy (anonymous):

how would you work this out? composite function f ◦ g f(x) = 2e^x g(x) = ln(x − 2) i worked it out : 2e^(ln(x-2) can u simplify this more?

OpenStudy (tkhunny):

Keep this in mind. \(\ln\left(e^{x}\right) = x\cdot\ln(e) = x\) This might lead you to the following conclusion... \(e^{\ln(x)} = x\)

OpenStudy (anonymous):

so is it 2(x-2) ?

OpenStudy (tkhunny):

Pretty good, but you will have to think about the Domain. It may have changed inadvertently along the way.

OpenStudy (anonymous):

so domain will be x cannot equal 2

OpenStudy (tkhunny):

The Domain of g(x) is \(x-2 \ge 0 \implies x \ge 2\) f(x) has no restriction.